North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6



Wednesday 26th April 2017

A bitterly cold snowy, sleety morning to start with, icy patches on the road from Teesside, but the sun came out later and, sheltered from the wind, it was reasonably warm - even better in the workshop with the heaters on! The regular Wednesday crew of Derek, Jon, and myself, with Ian in charge. John Furness from the NYMR was also in Deviation Shed first thing, sanding down the S15 tender and creating an enormous amount of dust such that the south end roller shutter doors had to be opened.

Fortunately he finished about 10.30.

Jon worked all day on carbon reduction around the steam chest area - there is still quite a lot more to remove. I continued to complete the valve cleaning after the excellent work done on Saturday and, with Ian's help, re-assembled both of them. The valves require the big nuts tightening and new cotter pins, but both are now sitting on the back of the work bench. After moving the tender back a yard or so to get a pit bridge in place, Derek set about cleaning the front around the draw bar area and drag box and has made a marked improvement. A plastic sheet was put over the water pump to keep the dirt out. Ian moved Gordon's electric box into the workshop, just on the left through the door near the arch formers, repositioned the 'Not to Be Moved' boards to hang from the beams above, put all the pieces of rope in a box next to the rag box, and had a general tidy up of the area.

Lunch was held in the A4 cabin as it was too cold for a picnic lunch outside. After lunch Ian started removing carbon from the left hand cylinder ports. A number of interested people round the shed today (including Nick Stringer), but didn't hear any bells ringing from the donation box!

 

Wednesday 19th April 2017

Yesterday at Grosmont a cool but fine day for Jon, Derek and myself with Ian in charge. After the necessary tea and HobNobs, Jon and Derek were back under the loco frames, cleaning the never ending muck and goo, while I got on with the NELPG Carbon Reduction Programme on the right hand valve as part of our green environmental initiative. This valve is far worse than the left one, with large chunks of carbon to remove, and, interestingly, so is the right hand exhaust port - signs of an atomiser problem? Ian finished cleaning the base of the snifting valve on the header and then, after lunch in the A4 cabin, started more carbon reduction on the left hand piston cover.

As we were getting quite a lot of visitors in the Shed, Ian and I made an executive decision to get rid of the old chair near the cage - it was filthy, broken and getting in the way of our collection box - so it's now in the skip. Apologies to anyone who still dearly loved it. We took time out to chat to our visitors and explain and show them what we were doing. Mention of the sums of money involved in keeping steam locomotives operating, and pointing out the donation box, managed to keep the bell ringing as the coins went in! Hope the Saturday team can keep up the good work. Bill Sharp and his wife also came to see us. Bill was measuring up the Q6 tender for his 5"gauge model.

On the Railway it was a full steam service with Pocket Rocket, the Standard Tank and the B1. All trains were well loaded, as would be expected with the school holidays still on. No timing for any move to the wheel drops as yet. Piglet is on a short holiday in Snowdonia for the rest of this week but back on Monday and still waiting for a meeting with Paul and Andy to agree the rest of the Q6 overhaul programme.

 

 

Saturday 22nd April 2017

Another day of the "Carbon Reduction Programme", well nearly, with Colin Foxton, Peter Ellis, James Swainston (first time at Deviation shed), Gordon Wells and Ian Pearson (Andy was away cycling round the Yorkshire Dales).

After a cup of tea and biscuits, Peter continued with removing carbon from the the right hand valve, while everyone else dismantled the staging round the smoke box and removed it to the NYMR boiler shop.

Colin then cleaned the right hand cylinder cover, Ian cleaned the left hand cover, and James worked on the steam chests where there was more carbon to remove. Both covers are now cleaned, lightly oiled and put under the seat bench with covers on top. Peter completed cleaning the right hand valve which is on the work bench to reassemble. Mark O'Brien was duty fitter and he said he had to re tap the threads in the fire box and measure up before ordering the new flue tubes - hopefully in the next few days.

Gordon started fitting some new sockets in the pit but they are not yet wired up. He also put a new bulb in the 110v indicator above the cage showing a red light and telling us the 110v system is on.

Quite a lot of visitors round the Shed on Saturday (including a visit by the SNG overhaul team, led by Darrin Crone and Martin Ashburner, on a social day out on the NYMR) with more money going in the collection box!

 

Wednesday 12th April 2017

A quiet day today with me opening up the workshop to be greeted by 2 packs of Hobnobs dangling from the ceiling, one dark and one milk chocolate. Lucky man! It was grey and cold, and poured with rain in the morning. No more fatalities to report.

Hob Nob

Derek Shorten joined me and, after a warming brew, set to work all day continuing with the cleaning of the wheels and frames outside and inside. It is a never ending job, but with some new scrapers, a lot more muck was removed. I continued cleaning the carbon off the L/H valve. Ian arrived about 11am, and met me at the signing on point as I had to leave for a meeting at Darlington. Ian continued cleaning the L/H valve until lunch time, and as it was a bit cool for the outside picnic area, Derek and Ian retreated to the A4 cabin. After completing the L/H valve Ian started on the R/H one: the ring and outer part are now clean with lots of hard carbon removed. However, there is still lots of carbon to remove from the rest of the valve.

Jon Bradley missed out on the Hobnobs this week, but at least he kept warm as he was driving the B1 on the 1030 ex Grosmont as it slipped past us on the way to Pickering.

The locomotive is not likely to get on the wheeldrop until the Green timetable kicks in after the Easter holiday, and there has been a meeting between Paul Middleton, Andy and Paul Hutchinson to agree the programme for the rest of the overhaul. The new flue tubes will be ordered after Easter, and advice is being sought on whether the elements should be renewed or are suitable for re-use.

No working party this weekend, but next Wednesday as normal. See you then - have a good Easter break.

 

Saturday 8th April 2017

We had a dedicated set of lads working on the Q6 today, Ian Pearson, Peter Ellis, Colin Foxton (he could only manage the morning due to other commitments, many thanks) Nic Ashmore who travelled up from Weymouth on the south coast on a motorbike, also Andy Lowes.

 

Nic Ashmore at the workbench - Andy Lowes

Both top eccentric rods were removed & safely stored after marking up. This was on the advice from Barney on Wednesday to remove them before moving the loco onto the wheel drops sometime in the future. We also rigged up an air line to the reverser & put the engine into mid gear. The expansion links were also tied up to stop them moving around.

 

Top eccentruc rids after removal - Andy Lowes

The new stuffing box for the regulator was fitted to the back head after rather a lot of cleaning of the studs & the face of the boiler where it mates up to. It was then removed & a gasket made, the stuffing box was then put into safe storage. This is because the regulator rod has to be pulled part way out to gain access into the boiler via the dome.

 

The stuffing box face after much cleaning - Andy Lowes

The short pipes to the cab gauges have been annealed see photo. Another attempt has been made at removing the plugs in the super heater header, good news one has moved a quarter of a turn on the right hand side!

 

Peter Ellis annealing lubrication pipes - Andy Lowes

Cleaning has continued on the rods & brake gear that has been removed from the loco. The anti vacuum valve that has been repaired by M-Machine had new gaskets made & new nuts fitted. The sealing ring that fits on the outside of the smoke box had the packing removed & the holes tapped out then new bolts fitted where required. This seals the anti vacuum valve to the smoke box.

As next Saturday (15th April) is Easter there will be no working party on the Q6 in Deviation shed.

The following Saturday (22nd April) Ian Pearson has agreed to be RO.

All Wednesdays will be as normal with Ian. We have lots of goodies in safe storage thanks to Nigel Hall for the donation. No excuse for not enjoying your time at Deviation shed with added attractions of trains going past the door, great fun!

Now for the list of jobs in no particular order:

  • Clean valves
  • Header plugs to remove
  • Clean loco & tender frames/wheels
  • Check & record wheel centres before taking out driving wheels
  • Once driving wheels are out, remove the axle boxes for examination & eccentric straps
  • Fit new axle box lubrication pipes (from stock above workshop) as leading & intermediate ones
  • Back head fittings, valves, water gauges etc to refurbish
  • Split pins on studs inside boiler to be replaced where missing or damaged
  • Regulator handle needs to be a good fit on the tapered square of the regulator rod, welding & dressing of the handle is required to achieve this
  • Full survey of the loco brake gear to find what requires repair or replacement
  • Thickness test required on both main steam pipes
 

Wednesday 5th April 2017

Double disaster yesterday - brewed up the tea to find the milk had gone off, and then discovered all the chocolate hobnobs had been eaten! Ian Pearson saved the day by going down to the Co-op and getting some more milk, but we - Derek Shorten, Jon Bradley and myself plus Ian - just had to put up with chocolate digestives!

In spite of this setback, I carried on with cleaning the left hand valve and chipping out the large amounts of hard carbon. Very stubborn to remove, and there is still more to clear from the left hand valve head. The right hand one has yet to be started. I also collected some new underkeep pads for the J72 from Armstrong Oilers which Fred had ordered, and which Derek agreed to take to Hopetown today.

Derek continued with cleaning the wheels and frames, while Jon and Ian removed and cleaned the brake tables and brake hangers. These are now all removed and cleaned and marked up, barring the trailing table which still needs cleaning. Nigel Hall arrived later in the morning to gloss paint the tender steps. Bill Dobson also called in during the afternoon to see how were getting on.

As it was such a nice day, it was decided to have lunch in the picnic area and watch the trains go by.

Because Ian had noticed that the eccentrics were hitting the eccentric rod, which is disconnected and fastened up, Barney, the MPD Senior Fitter, had a look at the situation and advised that we should remove the top two eccentric rods before we arrange any moves to the wheel drops.

 

 

Saturday 1st April 2017

Spring is in the air so we had a clean up in the shed, Bryan Orange with the JVs, Ian Pearson, Peter Ellis, Colin Foxton, Gordon Wells, Andy Lowes. Joan Lowes put a sick note in at last minute with a chesty cough.

A real good effort today the shed & workshop have both been cleaned out. The morning JVs cleaned the shed & stored the Q6 items off road 6 behind the bench to allow the railway to get their diesel shunter out to put the new engine into it. The M10 holes in the lower spark arrester were tapped out & the clamping plates bolted to it, also the missing plate was manufactured. The afternoon JVs concentrated on the workshop clean up, this involved all the tools on the wall being cleaned in kerosene & wiped with a cloth. The tool lockers were also tackled & what a difference it has made, we can now get tools out without getting dirty before we start a job. Please keep it this way by cleaning tools before returning them to their home, thanks!

The pile of bits from No29 in front of No5 have been moved up & put on pallets so No5 could be moved forward. This allowed our electrician with his helpful JVs to install the armoured cable into the pit for our new 110v sockets.

The bolts from the lower flange that hold the blast pipe in the smoke box were removed & the copper gasket put with the blast pipe. The six plugs three in each end of the super heater header do not want to come out, we have tried a bit of heat several times but they still remain firmly in the header.

The new stuffing box for the regulator was tried today, but the studs in the boiler back head require running down with a die nut to allow it to sit flush. Still quite a bit of preparation work is required to the boiler face to get this to fit properly, also a gasket to make.

So as you have seen lots of little fiddly part done preparation jobs for the future assembly of the loco today, this is what is known as ‘unseen work’ in the trade.

Now for Wednesday 5-4-17, there is almost 2 pints of milk in the fridge but no chocolate hobnobs!

  • Continue cleaning the valves
  • Grind anti vacuum valve in using fine paste only
  • Finish stripping loco brakes & clean
  • Clean wheels & under loco in preparation of taking the driving wheels out

 

 

Wrdnesday 29th March 2017

Ian reports quite a productive day with Nigel Bill, Derek Shorten, Jon Bradley and himself. I was absent this week waiting various deliveries at home (nuff said!) but Nigel was the real hero - would you believe he came to Grosmont on the train from Thornaby with all the school kids!

Jon kept back the chocolate Hobnobs until the 3'o'clocks (I have missed out again).

After the morning cup of tea, Jon and Nigel started work on removing brake gear. They took down the driving set first. This was cleaned and marked up and laid out on the timber frame along side the S15. After lunch the trailing set was removed but not cleaned as time ran out (Nigel had to get the 16 15 train back to Middlesbrough). Derek spent the day cleaning the driving set of wheels and surrounding frame work, while Ian worked all day removing more of the concrete from the smokebox to reveal the hidden nuts. He had to remove 2 barrow loads to gain access, but two on the right hand side and one on the left are now accessible. There is a little more of this super hard concrete to remove from under the rim to allow us to get at the other nuts.

At the same time John Furness was continuing his work on the Drewry shunter, rubbing down then painting. Bill Dobson came in to see how we were progressing, and there were also quite a few visitors looking round.

 

Saturday 25th March 2017

Crowds of people today, we parked in the national parks car park as requested due to a large green engine running with Terry Newman at the controls. I am not used to all these people crowding the platforms after three months with just NELPEG workers around Grosmont! It was nice to see a good appreciative public around our shed, it makes all the hard work we do on Q6 worthwhile.

So today we were four Ian Pearson (shed shunter driver commandeered) Alan Hardie, Peter (the plumber) Ellis & Andy Lowes.

The work today was to assemble the base of the spark arrestor onto the blast pipe & prepare the new bolts to hold the mesh in place. This was achieved apart from tapping the holes in the framework this will be done next Saturday if I remember to bring a tap along.

The centre right side rod was transported to the press in the running shed & both bushes were pressed out ready to be re-metalled & machined.

We had a very pleasant lunch break in hot sunshine at the picnic tables & Paul Middleton joined us to discuss where we are with the Q6. I told him we are ready to get the wheels out & it has been agreed to start with the drivers first, then return the loco to Deviation shed so freeing up the wheel drops for spring changes etc. This will not be this next week due to pre running season last minute jobs still to do, also Paul needs to agree costing for this work with NELPG before work starts on the second stage. First stage was removal of flue tubes with Mark O’Brian, now complete.

The anti vacuum valve has been cleaned & a crack was discovered in one of the four spokes of the casting. Neal Woods called in so I discussed with him the various methods of repairing it. Neal got in contact with our CME John Graham & will have a look at it on Monday evening at Hopetown, all very productive with quick results, I am well impressed thanks.

The side covers on the smoke box were also removed to expose the end plugs on the supper heater header. A dirty job but the dark hair took our plumber back to his youthful looks he said!

Lots of little fiddly preparation jobs today but good progress was made all the same.

Jobs for Wednesday 29th March:

  • Dig out more concrete around bolts for blast pipe & remove bolts
  • Loco brake gear to strip, clean & inspect
  • Continue stripping/cleaning valve assemblies
  • Clean loco ready to remove wheels
 

Wednesday 20th March

Derek Shorten and Ian Pearson were eventually joined by Jon Bradley and myself (too much chatting on the way through the MPD as well as Royal Scot to get round). The Q6 was not on the wheel drops and probably won't get there until Andy Lowes and Paul Hutchinson have discussed with Paul Middleton the next stage of the mechanical work, and the wheels/axle boxes in particular. No mice this week.

After a cup of tea and some non chocolate hobnobs, Jon and Derek set about cleaning frames all day. I got on with stripping down the left hand valve and cleaning off the hard carbon which was very stubborn to remove. There is still a lot of work to do on this valve, and the right hand one to make a start on. The valve has been loosely reassembled to keep the components in the right order and is back on the bench in the workshop.

Ian went to see Howard Smith at Armstrong Oilers to check out the item number for the Q6 underkeep pads, confirmed as '814B'. Back at the workshop he found we had 6 x '814B ' pads made at York, probably some fifty years ago, and 6 x '814B' pads made by Armstrong Oilers in 2014, giving 12 in all. The latter 6 pads have been put in tins and tubs of oil to soak, and are on the bench outside the workshop. There were also some J27 pads in the upstairs store - number '813B'

After lunch, in the warm in the Gresley cabin, Ian started to remove the anti vacuum valve from the header. It needed quite a bit of heat but, after an hour or so, with some assistance from Jon, it finally came free and is now on the bench in the workshop .

In the morning, it was good to see Bob Grey who called in for a chat with us all. Additionally, John Furness from the MPD spent the day sanding and painting the Drewry: the replacement engine is due to arrive next week.

 

 
  • «
  •  Start 
  •  Prev 
  •  1 
  •  2 
  •  3 
  •  4 
  •  5 
  •  6 
  •  Next 
  •  End 
  • »


Page 1 of 6